# Bayesian Optimization for Robust State Preparation in Quantum Many-Body Systems

### v1

@@ Abstract
With this most basic transfer-learning approach, we reach the highest fidelity achieved in the noisy environment with a fraction of the necessary number of runs in the disordered system. Comparing this protocol with the optimized result from the clean system tells us something about the changes in the cost landscape when disorder is included. The new optimum in the disordered system is very close in parameter space to that in the clean system, so we find disorder to only lead to a slight shift in the ideal strategy. This scenario is distinct from one realized in a previous study Mukherjee et al. (2020b) where the optimum location drastically changed with the onset of disorder. For a more detailed discussion of these two scenarios, see Appendix F which confirms our findings as a property of the system in question – not an artifact of optimization – via a brute-force search and examines a special case in our system that mirrors the scenario in the literature.
Short Ramps – Thus far, we have aimed to produce ramping protocols that achieve high fidelity both in an ideal as well as a noisy setting. The total ramp time $T$ has been kept fixed to that of the adiabatic experimental protocol. However, a major advantage of QOC techniques is that they can significantly reduce the total protocol time while retaining high fidelities Wigley et al. (2016); van Frank et al. (2016). To leverage this advantage, we run optimization with reduced final times $T=10\>\tau$, and $T=20\>\tau$, i.e., 10×10$\times$10 ×, or 5×5$\times$5 × shorter than the protocols previously considered. Based on our findings in the previous sections, we expect the global optimum of the optimization landscape to move only a little when introducing experimental disorder, so we chose to work in the ideal system.
@@
We consider the ramp resulting from the ramping time $T=20\>\tau$ to be optimal for an experimental implementation. While the fidelity in the disorderless scenario is still significantly lower than for optimized ramps with longer ramp times, we find the protocol to be extremely robust under the introduction of disorder. In fact, the fidelity at the strongest considered level of disorder $\sigma=0.2\>\hbar/\tau$ is significantly higher than for any other protocol considered in this work. Similar to previous studies van Frank et al. (2016); Mukherjee et al. (2020b), we attribute this result to the shorter protocol time presenting less of an opportunity for the accumulation of errors. This interpretation is supported by the significantly lower standard deviation of the fidelity in the noisy setting. At the experimental disorder level $\sigma=0.1\>\hbar/\tau$, while the mean fidelity of the $T=20\>\tau$ ramp is similar to that of the ramp optimized for $T=100\>\tau$, the standard deviation is lower by more than a factor of $4$.
Experimental Proposal – Based on the findings of the previous sections, we propose the following strategy and best practices for the realization of the BO approach in an experiment:
@@
(iii) Transfer learning: In case sensitivity analysis fails to recover the optimum, it is favorable to train on the experiment directly. In this scenario, we have demonstrated that making use of simulation data in a simple transfer-learning approach can significantly boost performance while reducing experimental costs.
@@
Discussion & Outlook – We applied a Bayesian optimization approach to optimize parameter ramps for state preparation. The optimization scheme was demonstrated to be efficient at finding protocols that outperform manual design not only in an ideal system but also in the presence of significant disorder – identifying a ramp that is both 5×5$\times$5 × faster as well as twice as close to ideal fidelity as the manual reference. The superior performance holds when the training data itself is noisy, e.g., when training directly on an experiment, but at an increased cost of training.
@@
We propose several approaches to decrease experimental costs in cases where numerical simulation of the system is available. Namely, we demonstrated that optimized solutions for the ideal system can be intrinsically robust due to a significant reduction in protocol duration and the surrogate model of the fidelity-landscape naturally provides information about the most sensitive parameters to tune in an experiment. When learning directly on the experiment, prior training in a simulated environment was shown to significantly reduce the total cost. This also gives hope for scenarios where simulation is available only in small systems, where transfer learning techniques may still be able to extract an advantage.
In more general terms, we believe a further reduction in the number of experimental runs by means of transfer learning and by improving the ramps’ parametrization to be the most pressing issues.
We expect the optimized protocols developed here to be instrumental in realizing larger FQH states – e.g., by merging several $4\times 4$ plaquettes into a larger system Palm et al. (2023) – and different filling factors. In particular, the substantial speedup and robustness offered by protocols optimized using machine learning can allow the preparation of these more complex quantum states within the already available coherence times of contemporary ultracold-atom experiments.
The more general optimization strategy, we believe to be widely applicable to boost experimental efficiency and improve the interface between classical and quantum simulation.
@@ Appendix A Bayesian Optimization
Bayesian optimization (BO) Shahriari et al. (2016); Garnett (2023) is an optimization method designed to find the optimum of a – possibly noisy and expensive to evaluate – black box function. In our case, the task is finding the optimal ramping protocol $\hat{H}(t)$ for the time-dependent Hamiltonian that transforms a given initial state |$\ket{\psi_{\mathrm{initial}}}$ into the target state |$\ket{\psi_{\mathrm{target}}}$. The quality of a given ramping protocol is quantified by the fidelity
@@
To find optimal values for the parameter set $\mathbf{x}=\{x_{j}\}_{j}$, a BO loop, based on the open-source scikit-optimize package, is run in the following way: First, the search is initialized by randomly choosing some $\mathbf{x}_{i}$ – typically $\mathcal{O}(10)$ points – and evaluating the corresponding cost function values $C(\mathbf{x}_{i})$ – either by running an experiment or, in our case, via numerical simulation. We then use Gaussian processes (GP) Rasmussen and Williams (2005) to build a surrogate model $\tilde{C}(\mathbf{x})$ of the cost function over the entire parameter space, based on the set of evaluated points. A GP is a generalization of the Gaussian distribution, defining a distribution over functions by replacing the mean vector and covariance matrix by the mean function $m(\mathbf{x})$, and the kernel $k(\mathbf{x},\mathbf{x^{\prime}})$ Rasmussen and Williams (2005). Given a set of previously evaluated points, the GP model predicts the value of the cost function over the entire parameter space and gives a measure of the model’s uncertainty in less-explored areas. Key properties of the GP model, such as the smoothness of and variation between sampled functions are controlled via a set of hyperparameters, which are continually adjusted during training via maximum-likelihood estimation. These hyperparameters include the noise level $\sigma_{F}$ of the data as well as the length scales $l$ which give the characteristic scale on which functions sampled from the GP change their value. Control over these parameters can be a valuable tool to achieve optimum BO performance. For example, manually setting $\sigma_{F}=0$ for training in the clean system can boost convergence and contribute to the low number of iterations needed compared to training in the noisy setting. We also demonstrate how limiting the lower bound of $l$ – constraining the maximum complexity of the constructed model – can improve robustness while speeding up convergence in Appendix G.
@@
The suggestion $\mathbf{x}_{i}$ is made by the acquisition function, which can prioritize either exploitation – choosing an $\mathbf{x}_{i}$ which minimizes $\tilde{C}(\mathbf{x})$, or exploration – choosing an $\mathbf{x}_{i}$ where $\tilde{C}(\mathbf{x})$ has high uncertainty. In the lower confidence bound (LCB) Srinivas et al. (2012) acquisition function
@@ Appendix B Exact Diagonalization Simulation
Exact-diagonalization (ED) simulations are used to perform ground-state search and time evolution for the Harper-Hofstadter Hamiltonian Harper (1955); Hofstadter (1976) on a $4\times 4$ lattice with $N=2$ bosons. The parameters $t_{x/y}$ and $\Delta_{x/y}$ are the control fields tunable in the experiment – representing tunneling amplitudes and global lattice tilts respectively. Site-dependent chemical potentials $\mu_{x,y}$ may be introduced to model disorder in the experiment.
For efficiency, we exploit the $\mathrm{U}(1)$ symmetry Zhang and Dong (2010), resulting from particle number conservation. As in the experiment, the initial state is prepared by manually placing the 2 bosonic particles in the bottom-left corner of the $4\times 4$ lattice, i.e.,
@@
Time evolution is performed using the Runge-Kutta method of order $\geq 4$ with a maximum step-size $\delta t_{\mathrm{max}}=0.1$.
@@
Appendix C Experimental Noise Model
The most important source of disorder in the experimental system are static chemical-potential offsets. We consider two separate contributions to this offset potential.
@@ Appendix C Experimental Noise Model
Firstly, we consider random offsets with a maximum strength of $\sim 10-15\mathrm{Hz}$ that are uncorrelated and follow the same distribution on each lattice site. Secondly, we incorporate a chemical-potential offset on the edge of the system caused by the digital micromirror devices (DMDs) projecting the walls of the lattice. The offset is modeled as having strength $2\mu_{\mathrm{DMD}}$ in the corners of the lattice and $\mu_{\mathrm{DMD}}$ everywhere else on the edges, equivalent to a harmonic potential affecting the entire $4\times 4$ lattice. The strength $\mu_{\mathrm{DMD}}$ of this type of disorder is taken to follow the same distribution as the random offsets, that is $\mu_{\mathrm{DMD}}$ may have either sign and the two types of disorder are similarly strong. As we expect the disorder to have mean 00 and maximum strength $\sim 10-15\mathrm{Hz}$ which is about a quarter of the hopping amplitude $\hbar/\tau\sim 40\>\mathrm{Hz}$ in the final FQH state, which sets the unit of energy, we model the disorder $\mu\ [\hbar/\tau]$ as being Gaussian-distributed with mean 00 and standard deviation $\sigma=0.1\>\hbar/\tau$. For each numerical time evolution, we sample $4\times 4+1$ independent values from this distribution – one for each single-site offset and one for the global harmonic one.
@@ Appendix G Making Use of the Trained Model
Based on the analysis presented in the main text, as well as the brute-force mapping of the cost landscape detailed above, we are confident that our system and disorder level realize scenario (i) – both for long ($T=100\>\tau$) and short ($T\leq 20\>\tau$) ramping protocols.
Figure 9: GP Model: The optimization landscape of the global tilts (hopping amplitudes) with the remaining 2 parameters fixed to their optimum value is shown in the left (right) plot. The red dot marks the location of the optimum identified by BO
However, dramatic changes in the global optimum’s location with the introduction of noise have been reported in the literature Mukherjee et al. (2020a) indicating scenario (ii). To look for (ii) in our system and test our conjecture, we turn to the intermediate ramp time $T=50\>\tau$. Here, in the clean system, BO identifies an optimum that it predicts to be extremely narrow in parameter space. The corresponding control path and GP model are illustrated in Fig. 9. The figure also shows the model’s uncertainty, which becomes high, even very close to the discovered optimum. To accommodate the narrow optimum, the length-scale hyperparameters of the GPs become very low during training, making the uncertainty away from evaluated points high.

### v2

@@ 5 Training with Noise
With this most basic transfer-learning approach, we reach the highest fidelity achieved in the noisy environment with a fraction of the necessary number of runs in the disordered system. Comparing this protocol with the optimized result from the clean system tells us something about the changes in the cost landscape when disorder is included. The new optimum in the disordered system is very close in parameter space to that in the clean system, so we find disorder to only lead to a slight shift in the ideal strategy. This scenario is distinct from one realized in a previous study [34] where the optimum location drastically changed with the onset of disorder. For a more detailed discussion of these two scenarios, see Appendix F which confirms our findings as a property of the system in question – not an artifact of optimization – via a brute-force search and examines a special case in our system that mirrors the scenario in the literature.
6 Short Ramps
Thus far, we have aimed to produce ramping protocols that achieve high fidelity both in an ideal as well as a noisy setting. The total ramp time $T$ has been kept fixed to that of the adiabatic experimental protocol. However, a major advantage of QOC techniques is that they can significantly reduce the total protocol time while retaining high fidelities [16, 10]. To leverage this advantage, we run optimization with reduced final times $T=10\>\tau$, and $T=20\>\tau$, i.e., 10×10$\times$10 ×, or 5×5$\times$5 × shorter than the protocols previously considered. Based on our findings in the previous sections, we expect the global optimum of the optimization landscape to move only a little when introducing experimental disorder, so we chose to work in the ideal system.
@@ 6 Short Ramps
We consider the ramp resulting from the ramping time $T=20\>\tau$ to be optimal for an experimental implementation. While the fidelity in the disorderless scenario is still significantly lower than for optimized ramps with longer ramp times, we find the protocol to be extremely robust under the introduction of disorder. In fact, the fidelity at the strongest considered level of disorder $\sigma=0.2\>\hbar/\tau$ is significantly higher than for any other protocol considered in this work. Similar to previous studies [10, 34], we attribute this result to the shorter protocol time presenting less of an opportunity for the accumulation of errors. This interpretation is supported by the significantly lower standard deviation of the fidelity in the noisy setting. At the experimental disorder level $\sigma=0.1\>\hbar/\tau$, while the mean fidelity of the $T=20\>\tau$ ramp is similar to that of the ramp optimized for $T=100\>\tau$, the standard deviation is lower by more than a factor of $4$.
7 Experimental Proposal
Based on the findings of the previous sections, we propose the following strategy and best practices for the realization of the BO approach in an experiment:
@@ 7 Experimental Proposal
Direct implementation of a numerically optimized (local adiabatic) control path has previously been used successfully to realize few-particle crystallization in an experiment [35].
@@
(iii) Transfer learning: In case sensitivity analysis fails to recover the optimum, or the system is intractable numerically, it becomes favorable to train on the experiment directly. In Section 5, we have demonstrated that making use of simulation data in a simple transfer-learning approach can significantly boost performance while reducing experimental costs (represented by the number of simulation runs in the disordered system). In Section H we also find that finding high-fidelity state-preparation protocols for a $4$-particle Pfaffian-like state [36] on a larger $L_{x}=L_{y}=6$ lattice is made much easier by transferring the results for the smaller state obtained in Section 4. This presents an avenue that could extend efficiency gains to experimental systems beyond the reach of classical simulation.
8 Discussion and Outlook
@@
We applied a Bayesian optimization approach to optimize parameter ramps for state preparation. The optimization scheme was demonstrated to be efficient at finding protocols that outperform manual design not only in an ideal system but also in the presence of significant disorder – identifying a ramp that is both 5×5$\times$5 × faster as well as twice as close to ideal fidelity as the manual reference. The superior performance holds when the training data itself is noisy, e.g., when training directly on an experiment, but at an increased cost of training.
@@
We propose several approaches to decrease experimental costs in cases where numerical simulation of the system is available. Namely, we demonstrated that optimized solutions for the ideal system can be intrinsically robust due to a significant reduction in protocol duration and the surrogate model of the fidelity-landscape naturally provides information about the most sensitive parameters to tune in an experiment. The reductions in experimental cost achievable by leveraging numerical resources are especially valuable to systems of ultracold atoms in optical lattices, where estimating the figure of merit usually requires many experimental snapshots.
Even when learning directly on the disordered system representing an experiment, prior training in a simulated environment was demonstrated to significantly reduce the total cost in Section 5.
The central challenge ahead lies in extending efficient optimization to system sizes that are not tractable using numerical simulation. Promising recent results in this regard were obtained by Y.-J. Xie et al. [20], who report efficient transfer of BO to iteratively larger system sizes in a Heisenberg chain. For the FQH system discussed here, we demonstrate good transferability to a larger $4$-particle state with different model parameters in Section H. In particular, this marks an example, where the substantial speedup offered by protocols optimized using machine learning can allow the preparation of a more complex quantum state within the already available coherence times of contemporary ultracold-atom experiments.
By using either such a transfer-learning approach or alternative schemes merging several FQH plaquettes [37], we expect the optimized protocols developed here to be instrumental in realizing larger FQH states, such as the Pfaffian state, and different filling factors. The more general optimization strategy, we believe to be widely applicable to boost experimental efficiency and improve the interface between classical and quantum simulation.
Data and Code Availability – The data presented in the figures is available at https://github.com/TizianBlatz/bo-state-preparation, which also contains a minimal working example of the optimization routine.
@@ Appendix A Bayesian Optimization
Bayesian optimization (BO) [14, 15] is an optimization method designed to find the optimum of a – possibly noisy and expensive to evaluate – black box function. In our case, the task is finding the optimal ramping protocol $\hat{H}(t)$ for the time-dependent Hamiltonian that transforms a given initial state |$\ket{\psi_{\mathrm{initial}}}$ into the target state |$\ket{\psi_{\mathrm{target}}}$. The quality of a given ramping protocol is quantified by the fidelity
@@
To find optimal values for the parameter set $\mathbf{x}=\{x_{j}\}_{j}$, a BO loop, based on the open-source scikit-optimize package, is run in the following way: First, the search is initialized by randomly choosing some $\mathbf{x}_{i}$ – typically $\mathcal{O}(10)$ points – and evaluating the corresponding cost function values $C(\mathbf{x}_{i})$ – either by running an experiment or, in our case, via numerical simulation. We then use Gaussian processes (GP) [29] to build a surrogate model $\tilde{C}(\mathbf{x})$ of the cost function over the entire parameter space, based on the set of evaluated points. A GP is a generalization of the Gaussian distribution, defining a distribution over functions by replacing the mean vector and covariance matrix by the mean function $m(\mathbf{x})$, and the kernel $k(\mathbf{x},\mathbf{x^{\prime}})$ [29]. Given a set of previously evaluated points, the GP model predicts the value of the cost function over the entire parameter space and gives a measure of the model’s uncertainty in less-explored areas. Key properties of the GP model, such as the smoothness of and variation between sampled functions are controlled via a set of hyperparameters, which are continually adjusted during training via maximum-likelihood estimation. These hyperparameters include the noise level $\sigma_{F}$ of the data as well as the length scales $l$ which give the characteristic scale on which functions sampled from the GP change their value. Control over these parameters can be a valuable tool to achieve optimum BO performance. For example, manually setting $\sigma_{F}=0$ for training in the clean system can boost convergence and contribute to the low number of iterations needed compared to training in the noisy setting. We also demonstrate how limiting the lower bound of $l$ – constraining the maximum complexity of the constructed model – can improve robustness while speeding up convergence in Appendix G.
@@
The suggestion $\mathbf{x}_{i}$ is made by the acquisition function, which can prioritize either exploitation – choosing an $\mathbf{x}_{i}$ which minimizes $\tilde{C}(\mathbf{x})$, or exploration – choosing an $\mathbf{x}_{i}$ where $\tilde{C}(\mathbf{x})$ has high uncertainty. In the lower confidence bound (LCB) [38] acquisition function
@@ Appendix B Exact Diagonalization Simulation
Exact-diagonalization (ED) simulations are used to perform ground-state search and time evolution for the Harper-Hofstadter Hamiltonian [31, 32] on a $4\times 4$ lattice with $N=2$ bosons. The parameters $t_{x/y}$ and $\Delta_{x/y}$ are the control fields tunable in the experiment – representing tunneling amplitudes and global lattice tilts respectively. Site-dependent chemical potentials $\mu_{x,y}$ may be introduced to model disorder in the experiment.
For efficiency, we exploit the $\mathrm{U}(1)$ symmetry [39], resulting from particle number conservation. As in the experiment, the initial state is prepared by manually placing the 2 bosonic particles in the bottom-left corner of the $4\times 4$ lattice, i.e.,
@@
Time evolution is performed using the Runge-Kutta method RK5(4) [40] with a maximum step-size $\delta t_{\mathrm{max}}=0.1$.
@@
Appendix C Experimental Disorder Model
For a more realistic description of the experimental system, we consider disorder in the shape of static chemical-potential offsets that can be divided into two components: One as random offsets, and another as a harmonic confinement.
The random offsets arise due to light scattering on surfaces, introducing disorder in the optical lattice. We model these offsets as uncorrelated with a maximum strength of $\sim 10-15\>\mathrm{Hz}$, or about a quarter of the tunneling amplitude $\hbar/\tau\sim 40\>\mathrm{Hz}$ in the final FQH state which sets our unit of energy. The harmonic confinement is introduced by the optical potentials, or walls, that defines the region of interest. The walls are designed to have a Gaussian shape and are holographically projected by a digital micromirror device (DMD) [41]. The DMD is located in the Fourier plane, and the finite aperture of the DMD limits the sharpness of the walls, resulting in edge offsets that may be modeled as a harmonic confinement.
@@ Appendix C Experimental Disorder Model
Thus, on the $4\times 4$ lattice, the offset is modeled as having strength $2\mu_{\mathrm{DMD}}$ in the corners of the lattice and $\mu_{\mathrm{DMD}}$ everywhere else on the edges. The strength $\mu_{\mathrm{DMD}}$ of this type of disorder is taken to follow the same distribution as the random offsets, that is $\mu_{\mathrm{DMD}}$ may have either sign and the two types of disorder are similarly strong.
In the simulation, we model the disorder $\mu\ [\hbar/\tau]$ as being Gaussian-distributed with mean 00 and consider a standard deviation of $\sigma=0.1\>\hbar/\tau$ to be the experimentally realistic disorder level. For each numerical time evolution, we sample $4\times 4+1$ independent values from this distribution – one for each single-site offset and one for the global harmonic one.
We choose this disorder model as opposed to some other common experimental imperfections such as elevated temperatures and particle losses, as these can be well controlled in the small system by postselecting runs with the correct low-energy initial particle configuration and final particle number via density snapshots.
@@ Appendix G Making Use of the Trained Model
Figure 9: Different GP Models: Results of $2$ independent BO runs with $500$ iterations each for an intermediate ramp time $T=50\>\tau$. Based on different hyperparameter settings, the models identify different optima. Each subfigure ((a), and (b)) shows the optimized ramping protocol (left) and the modeled fidelity-landscape of the $8$ control parameters around the optimum (right). (a) The model after training with the same parameter settings as in the main text. An optimum with fidelity $F=81\%$ is identified. The optimum is very narrow in parameter space and is especially sensitive to the $\Delta_{y}$ control points. The model uncertainties (shaded areas) away from the optimum are large. (b) GP model with a 10×10$\times$10 × lower bound for the length scale hyperparameter than in (a). The lower-complexity model identifies a different optimum in the optimization landscape that lies lower in terms of fidelity ($F=70\%$) but is much wider. The model uncertainty around the optimum is very low.
@@
Based on the analysis presented in the main text, as well as the brute-force mapping of the cost landscape detailed above, we are confident that our system and disorder level realize scenario (i) – both for long ($T=100\>\tau$) and short ($T\leq 20\>\tau$) ramping protocols. However, dramatic changes in the global optimum’s location with the introduction of noise have been reported in the literature [19] indicating scenario (ii). To look for (ii) in our system and test our conjecture, we turn to the intermediate ramp time $T=50\>\tau$. Here, in the clean system, BO identifies an optimum that it predicts to be extremely narrow in parameter space. The corresponding control path and GP model are illustrated in Fig. 9. The figure also shows the model’s uncertainty, which becomes high, even very close to the discovered optimum. To accommodate the narrow optimum, the length-scale hyperparameters of the GPs become very low during training, making the uncertainty away from evaluated points high.
@@
Appendix H Larger System
As an outlook beyond the minimal two-particle Laughlin-like state at filling factor $\nu=1/2$ that has been realized experimentally, we conclude with first results on the $N=4$, $L_{x}=L_{y}=6$ Pfaffian-like state at $\nu=1$ proposed by F. Palm et al. [36]. Following their work, we use the system parameters $U=4\>\hbar/\tau$ and $\phi/2\pi=0.16$. We then optimize the ground-state preparation from an initial state of particles pinned in the corner of the system in a $1\times 4$ configuration.
Figure 11: Transfer to Larger System: BO training for the $N=4$, $L_{x}=L_{y}=6$, $\nu=1$ state. (a) The optimized protocol for the two-particle $\nu=1/2$ state from Section 4. (b) Optimized ramping protocol for the larger $\nu=1$ state obtained by transferring training data: $50$ protocols from the training of (a) are used to initialize the surrogate model. (c) Comparison between training initialized with $50$ transferred points and random initialization with $10$ points. The transfer approach is much more efficient, reaching $F>70\%$ in less than $200$ iterations. The shaded orange region marks the evaluation of transferred data points. The optimized Laughlin ramp prepares the larger $\nu=1$ state with $F=13\%$ ($\mathrm{Iteration}=0$; orange) – significantly better than random protocols, but far from optimal.
While pinning the particles in a corner of the system becomes increasingly restrictive when going to larger systems, we choose this state to most naturally extend the optimization problem from the $N=2$ case. Combined with the larger Hilbert-space dimension $\mathrm{dim}(\mathcal{H}^{N=4})=82251$, compared to $\mathrm{dim}(\mathcal{H}^{N=2})=136$, this makes for a significantly more challenging preparation task, where randomly chosen protocols are expected to yield almost vanishing fidelity. As the goal of this work is to allow the preparation of more complex states using contemporary experimental setups, we perform optimization with the same total protocol duration and restrictions imposed by Floquet heating as previously for the smaller system. That is, we again optimize protocols of total duration $T=100\>\tau$ with the coupling in x-direction only active for a shorter time $T_{x}=40\>\tau$. We wish to answer two main questions:
  1. (i)
    Can we still find good preparation protocols for the larger system? How do these compare to the adiabatic protocol for the smaller system?
  2. (ii)
    Can we use the knowledge gained in the small system to boost optimization efficiency for the more complex state?
In order to answer these questions, we perform two separate BO runs with $500$ iterations each. The first run follows the procedure described in Section 4, using $10$ random initial points, while the second run uses initial points collected from training on the small systems following the initialization algorithm described in Section E. The performance of these two approaches is compared in Fig. 11. While the ‘cold-start’ approach that uses random initialization continues to find better and better protocols as training goes on, the rate of improvement is low, reaching only $F=22\%$ after $500$ iterations. The transfer approach performs much better, reaching fidelities as high as $F=75\%$. While the optimized protocol for the $N=2$ system is not directly applicable to the larger system ($F=13\%)$, the qualitative strategy does not change. Fig. 12 compares the evolution of the density profile for the optimized protocol in the two system sizes.
Figure 12: Particle Density: Evolution of the density profile from the initial state of pinned particles comparing different system sizes. The plots on the left show the density distribution at different times during the ramp. The rightmost plot shows the density profile of the (target) ground state. (a) Smaller system of $2$ particles on $4\times 4$ sites evolving according to the optimized protocol from Section 4. (b) Larger system of $4$ particles on $6\times 6$ sites evolving according to the optimized protocol obtained via transfer learning.
[…]